OpenStudy (anonymous):

given two points M and N on the coordinate plane, find the slope of line MN, and state the slope of the line perpendicular to line MN: M(9,6) N(1,4)

OpenStudy (anonymous):

The answer of the slope is 4 :) here's the solution: \[Slope formula: x _{2}-x _{1}/y _{2}-y _{1} \] 

OpenStudy (anonymous):

Then Substitute: \[\frac{ 1-9 }{ 4-6 } = \frac{ -8 }{ -2 }= 4\] that's how i get the 4 :)